OpenStudy (anonymous):

giving medals One number is 4 more than another, and their sum is 60. What is the smaller number? If x = the larger number and y = the smaller number, then which of the following systems could be used to solve the problem? x + y = 60 and -x + y = 4 x + y = 60 and x - y = 4 x + y = 60 and x + y = 4

OpenStudy (anonymous):

a b c

OpenStudy (ajprincess):

it is given that one number is 4 more than the other. so the larger number must be 4 more than the smaller number. so we get y+4=x their sum is given is 60. ie the addition of those two numbers give 60. x+y=60

OpenStudy (anonymous):

ok the second one

OpenStudy (anonymous):

or the third

OpenStudy (anonymous):

hhmmmmmmmmmmmmm

OpenStudy (ajprincess):

y+4=x the diffrence between the numbers is 4. ie the subtraction of smaller number from the larger number gives 4

OpenStudy (anonymous):

yup the second one

OpenStudy (ajprincess):

yup:)